# Improving Data Curation of Software Vulnerability Patches through Uncertainty Quantification

### v1

@@ Abstract
In this paper, we propose an approach employing Uncertainty Quantification (UQ) to curate datasets of publicly-available software vulnerability patches. Our methodology leverages machine learning models that incorporate UQ to differentiate between patches based on their potential utility. We begin by evaluating a number of popular UQ techniques, including Vanilla, Monte Carlo Dropout, and Model Ensemble, as well as homoscedastic and heteroscedastic models of noise. Our findings indicate that Model Ensemble and heteroscedastic models are the best choices for vulnerability patch datasets. Based on these UQ modeling choices, we propose a heuristic that uses UQ to filter out lower quality instances and select instances with high utility value from the vulnerability dataset. Using our approach, we observe an improvement in predictive performance and significant reduction of model training time (i.e., energy consumption) for a state-of-the-art vulnerability prediction model.
Index Terms:
Software Vulnerability, Vulnerability Patch, Data Quality, Machine Learning
@@ I Introduction
To close this gap, research groups have curated historic vulnerability databases using: 1) manual validation of each patch for quality; 2) automated approaches based on data selection heuristics; or 3) machine learning-based approaches trained on small manually-curated datasets. These projects face two primary challenges. First, they rely on publicly disclosed vulnerability information, such as from the NVD, which recent research shows is often missing important information or inaccurate [13, 14]. For example, Tan et al. found that out of 6,628 CVEs, only 66.57% contained references to patches, and of these, 32.79% were incorrect [13]. Heuristics and machine learning models do not eliminate the errors present in these datasets and could even exacerbate them. Second, manual review of vulnerability patches to ensure high quality results in very limited dataset sizes. This effort is further complicated by the need for some familiarity with different software projects to validate each patch. Despite these costly manual efforts, the data can still suffer from significant quality issues [15].
@@
Aside from these, prior works mostly focus on patch data quality when the patches are curated or removed from an existing dataset [15], although the improved patch datasets are often evaluated via intended down-stream applications [13, 14]. We argue that data quality and data usefulness (i.e., the data utility value) are two related but distinct concepts and the data quality and the data usefulness span in a spectrum. Prior works often lack of systematic or algorithmic methods to assess patch quality and usefulness.
@@
To address this question, we compare various UQ techniques combining two data distribution modeling approaches, homoscedastic and heteroscedastic [17, 21], with three UQ estimation methods: Vanilla, Monte Carlo Dropout, and Model Ensemble [16, 18, 19]. Our evaluation indicates that prediction confidence from the Vanilla approach alone is insufficient as a UQ measure. Additional uncertainty measures from Monte Carlo Dropout and Model Ensemble reveal higher uncertainty as the quality of vulnerability patches declines. Moreover, UQ models incorporating noise distributions, such as heteroscedastic, respond better to changes in data quality and usefulness in a software vulnerability dataset. An UQ model that disentangles epistemic and aleatoric uncertainties is particular useful. Epistemic uncertainty can serve as a proxy for the usefulness of security patches while aleatoric uncertainty the quality or the noise level of the patches.
TABLE I: Vulnerability patch datasets.
Dataset Characteristics
language # projects # patches
SecBench [22] $>1$ 114 676
@@
VulCurator [32] Python 1 290
VulnCode-DB [33] $>1$ 3681
@@ II Aleatoric and Epistemic Uncertainty in Software Vulnerability Patch Data Curation
As exhibited in Table I, curating software vulnerability patch data has been attempted numerous times. Ensuring the high quality of vulnerability patch datasets requires a significant amount of manual effort in curating patches. For example, the SAP dataset [31] contains 1,282 patches manually curated by teams of developers at SAP through their daily work. Such manual data curation approaches face two main challenges. Firstly, they are difficult to scale due to the extensive manual inspection required and the limited availability of software security expertise. Secondly, these studies and datasets often arise from identifying quality issues in previous datasets, resembling a game of ”whack-a-mole” where new quality issues continue to emerge despite exhaustive and sometimes heroic efforts. Recently, machine learning-based approaches have emerged. In ML-based approaches, researchers manually curate an initial high-quality set of ”seed” vulnerability patches and use them to train a machine learning model, which then automatically identifies vulnerability patches [36, 37, 38, 3, 39, 4, 13, 14]. Figure 1 illustrates these two approaches.
@@
Epistemic uncertainty is only present in ML-based approaches and arises from the model’s lack of knowledge about the problem. It can manifest as variance in the model’s predictions using the same training dataset. For example, different training runs of a deep neural network may yield different model weights, leading to varied predictions, which illustrates epistemic uncertainty. It is anticipated that adding more training data can eventually constrain the model weights, thereby reducing epistemic uncertainty [40, 41, 42, 43]. However, this process can become computationally expensive if more training data instances are added indiscriminately, and there is no assurance of improved predictive performance if the added training data instances are of low quality.
In this paper, we propose an ML-based software vulnerability patch data curation process that addresses quality issues by measuring both epistemic and aleatoric uncertainties. This dual uncertainty measurement, along with our EHAL heuristic, defined in Section IV, aims to enhance the efficacy of ML-based vulnerability data curation.
@@
There are numerous approaches for quantifying epistemic and aleatoric uncertainty. We limit our focus to machine learning-based patch data curation models that are based on neural networks. To answer the question of which UQ approaches are suitable for the specific domain of vulnerability patch datasets, we first introduce three UQ approximation techniques in Section III-A. Following that, Section III-B discusses two different assumptions underpinning the representation of vulnerability data, specifically whether it is homoscedastic or heteroscedastic. Subsequently, we examine metrics that can effectively gauge the accuracy of the measured uncertainty in Section III-C. Finally, we discuss evaluation experiment setup and the results in Section III-D and Section III-E, respectively.
@@ III-A2 Monte Carlo Dropout
Monte Carlo Dropout is a popular regularization technique used in training neural networks, where a neural network layer’s output is connected to a dropout layer. With it, we drop, i.e, zero out, the activation of the connected neurons at probability $P_{d}$ during training. When dropouts are used to for UQ, we also turn on dropout during inference time, which provides a sampling of a single set of neural network weights [16]. In essence, with Monte Carlo Dropout, we approximate probability distribution $p(y|x,D,\theta)$ with a discrete probability distribution, i.e., $p(y|x,D,\theta_{i})$, $i=0,1,2,\ldots,T-1$ by sampling multiple sets of neural network weights with $T$ stochastic passes.
@@ III-C Uncertainty Measures
In the context of homoscedastic models, according to the Law of Total Variance, literature typically refers to the entropy defined in equation (1) as a measure for total uncertainty, the quality defined in equation (III-C) as a measure for data or aleatoric uncertainty, and the mutual information estimated in equation (2) as a measure for model or epistemic uncertainty [51, 52, 53]. In the context of heteroscedastic models, also according to the Law of Total Variance, we obtain two probabilistic distributions: one representing aleatoric uncertainty, $p_{ale}(y|x)$, estimated from the expectation of $\sigma$, and the other representing epistemic uncertainty, $p_{epi}(y|x)$, computed using the variance of $\mu$ (see Figure 3). With these two distributions, we can compute two entropy values, denoted as $\mathcal{H}_{ale}(y|x)$ and $\mathcal{H}_{epi}(y|x)$ according to equation (1), which correspond to aleatoric and epistemic uncertainties respectively.
@@ III-D2 Datasets and Feature Representation
III-D3 Evaluation Procedure
To answer RQ1, we design and carry out a series of numerical experiments. For each experiment, we follow standard training, validation and testing protocols. Unless otherwise stated, the ratio of training and test datasets is $0.8:0.2$. From the training dataset, 10% is set aside as a validation dataset. During training, the majority class, i.e., non-vulnerability patches, are under-sampled to have a balanced training dataset, while during evaluation the test dataset is kept as is. The validation dataset is used to tune model hyperparameters, such as neural network size and training epochs. We use the Adam optimizer during training and select the best performing checkpoint based on the negative likelihood loss of the final model computed on the validation dataset. For homoscedastic models, we train with the negative log likelihood loss function (i.e., mathematically, the cross-entropy loss in two-class classification), and for heteroscedastic models, we train with a stochastic negative log likelihood loss function introduced by Kendall and Gal [17].
@@ III-E2 Separating Epistemic from Aleatoric Uncertainty
In this section, we demonstrate that epistemic and aleatoric uncertainties behave as expected when curating patch data. In Table III, we set aside 80% of patches as training data, train the patch identification model with increasing training dataset sizes (60%, 80%, and 100% of the training data), and then estimate the epistemic and aleatoric uncertainties on the test dataset (the remaining 20% of the patches). The table shows that as we increase the training data size, there is a significant reduction in epistemic uncertainty, i.e., a 1.5% to 9.3% decrease in epistemic uncertainty when the training data increased from 60% to 80% and then to 100%. In contrast, the aleatoric uncertainties saw smaller changes. Since an increase in the training data size indicates an improvement in the model’s knowledge about the problem space, which is captured by a reduction in epistemic uncertainty, it follows that in for data curation of software vulnerability patches, a UQ technique that disentangles epistemic and aleatoric uncertainties is required.
@@
Data ¯$\overline{\mathcal{H}}_{epi}$ $\Delta_{epi}\%$$\overline{\mathcal{H}}_{ale}$ $\Delta_{ale}\%$ 60% 0.794 0.0% 0.817 0.0% 80% 0.782 $\frac{0.794-0.782}{0.794}\approx 1.5\%$ 0.819 $\frac{0.817-0.819}{0.817}\approx-0.2\%$ 100% 0.709 $\frac{0.782-0.709}{0.782}\approx 9.3\%$ 0.790 $\frac{0.819-0.790}{0.819}\approx 3.5\%$
@@
Finally, we pose the question of whether modeling the noise in software patch data as heteroscedastic offer any advantage over modeling it as homoscedastic. The results in Table IV are obtained by averaging over all data quality shift intensities (from 0 to 0.4 with a step size of 0.1). The results for the heteroscedastic model are from using the Model Ensemble UQ approximation approach. The results show that the heteroscedastic model offers a significant advantage in predictive performance (higher F1-score) and superior or similar UQ quality (lower or nearly equal Brier Score). Given this observation, we conclude that for vulnerability patches, the UQ measures offered by the heteroscedastic models are more accurate than the homoscedastic models, due likely to the fact that the data distribution in vulnerability patches can vary between projects (and within a project over time), which is more accurately captured by heteroscedastic models.
@@
{mdframed}
  1. 1.
    From the three UQ models we considered, Model Ensemble produces the most accurate UQ estimation in response to artificial data quality degradation.
  2. 2.
@@
    To effectively assess the knowledge gained by a machine learning-based automated patch curation, it is crucial to separate epistemic and aleatoric uncertainty. Epistemic and aleatoric uncertainty are distinguishable during automated patch data curation.
  3. 3.
    Heteroscedastic models provide more accurate UQ estimation for patch data curation.
Input: $\mathbb{D}_{seed}$: Seed patch dataset
    $\mathbb{X}_{pool}$: Pool of patches to select from
    $n$: $n$ patches to curate from the pool
Output: $\mathbf{D}_{new}$: List of selected patches

### v2

@@ Abstract
In this paper, we propose an approach employing Uncertainty Quantification (UQ) to curate datasets of publicly-available software vulnerability patches. Our methodology leverages machine learning models that incorporate UQ to differentiate between patches based on their potential utility. We begin by evaluating a number of popular UQ techniques, including Vanilla, Monte Carlo Dropout, and Model Ensemble, as well as homoscedastic and heteroscedastic models of noise. Our findings indicate that Model Ensemble and heteroscedastic models are the best choices for vulnerability patch datasets. Based on these UQ modeling choices, we propose a heuristic that uses UQ to filter out lower quality instances and select instances with high utility value from the vulnerability dataset. Using our approach, we observe an improvement in predictive performance and a significant reduction of model training time (i.e., energy consumption) for a state-of-the-art vulnerability prediction model.
@@ I Introduction
To close this gap, research groups have curated historic vulnerability databases using: 1) manual validation of each patch for quality; 2) automated approaches based on data selection heuristics; or 3) machine learning-based approaches trained on small manually-curated datasets. These projects face two primary challenges. First, they rely on publicly disclosed vulnerability information, such as that from the NVD, which recent research shows is often missing important information or inaccurate [13, 14]. For example, Tan et al. found that out of 6,628 CVEs, only 66.57% contained references to patches, and of these, 32.79% were incorrect [13]. Heuristics and machine learning models do not eliminate the errors present in these datasets and could even exacerbate them. Second, manual review of vulnerability patches to ensure high quality results in very limited dataset sizes. This effort is further complicated by the need for some familiarity with different software projects to validate each patch. Despite these costly manual efforts, the data can still suffer from significant quality issues [15].
@@
Aside from these, prior works mostly focus on patch data quality when the patches are curated or removed from an existing dataset [15], although the improved patch datasets are often evaluated via intended down-stream applications [13, 14]. We argue that data quality and data usefulness (i.e., the data utility value) are two related but distinct concepts and the data quality and the data usefulness exist in a spectrum. Prior works often lack of systematic or algorithmic methods to assess patch quality and usefulness. Furthermore, existing efforts typically focus on dataset cleaning, aiming to address technical errors and inconsistencies. As we demonstrate, our data curation approach instead selects data points that maximize utility, ensuring that the resulting dataset is optimized for improving model performance and training efficiency.
@@
To address this question, we compare various UQ techniques combining two data distribution modeling approaches, homoscedastic and heteroscedastic [17, 21], with three UQ estimation methods: Vanilla, Monte Carlo Dropout, and Model Ensemble [16, 18, 19]. Our evaluation indicates that prediction confidence from the Vanilla approach alone is insufficient as a UQ measure. Additional uncertainty measures from Monte Carlo Dropout and Model Ensemble reveal higher uncertainty as the quality of vulnerability patches declines. Moreover, UQ models incorporating noise distributions, such as heteroscedastic, respond better to changes in data quality and usefulness in a software vulnerability dataset. An UQ model that disentangles epistemic and aleatoric uncertainties is particularly useful. Epistemic uncertainty can serve as a proxy for the usefulness of security patches while aleatoric uncertainty the quality or the noise level of the patches.
TABLE I: Vulnerability patch datasets organized according to collection approach.
Dataset Characteristics Collection Appr.
lang. # proj. # patch H1 H2 M
SecBench [22] >1>1 114 676
@@
VCMatch [14] C/C++ 10 1669
H1: The CVE-IDs are mentioned in a commit message in the project repository. H2: A URL to the patch is available in the NVD listing for the CVE. M: Patches are manually validated.
@@ II Aleatoric and Epistemic Uncertainty in Software Vulnerability Patch Data Curation
As shown in Table I, numerous efforts have been made to curate software vulnerability patch data. The primary approaches rely on heuristics like searching for CVE-IDs in commit messages and examining URLs provided in the NVD listing for CVEs (i.e., H1 and H2 in Table I), which are inherently noisy and incomplete. For instance, many links gathered through H2 do not point to the actual patch but rather to related code, while CVE-IDs in commits for H1 are infrequent and can be part of tangled commits that include changes unrelated to patching the vulnerability. Ensuring the high quality of vulnerability patch datasets, therefore, requires significant manual effort. This is reflected in the relatively small size of the manually validated datasets in Table I.For example, the SAP dataset [33] contains 1,282 patches manually curated by teams of developers at SAP through their daily work. Such manual data curation approaches face two main challenges. Firstly, they are difficult to scale due to the extensive manual inspection required and the limited availability of software security expertise. Secondly, they cannot address the issue of missing or broken links between CVEs and patches in the NVD.
Recently, machine learning-based approaches have emerged. In ML-based approaches, researchers manually curate an initial high-quality set of “seed” vulnerability patches and use them to train a machine learning model, which then automatically identifies vulnerability patches [34, 35, 36, 3, 37, 4, 13, 14]. Thus, ML-based approaches reduce the manual effort to curate larger datasets; however, they still suffer from two shortcomings: whether the instances curated are of low-quality or whether the instances contribute to improving down-stream applications, such as vulnerability prediction. Motivated by recent advances in uncertainty quantification in ML, we propose a data curation framework that addresses these two shortcomings as illustrated in Figure 1, which depicts the conceptual framework of the proposed approach.
@@
Epistemic uncertainty is only present in ML-based approaches and arises from the model’s lack of knowledge about the problem. It can manifest as variance in the model’s predictions using the same training dataset. For example, different training runs of a deep neural network may yield different model weights, leading to varied predictions, which illustrates epistemic uncertainty. It is anticipated that adding more training data can eventually constrain the model weights, thereby reducing epistemic uncertainty [40, 41, 42, 43]. However, this process can become computationally expensive if more training data instances are added indiscriminately, and there is no assurance of improved predictive performance if the added training data instances are of low quality or do not provide meaningful information to enhance the model.
Thus, the proposed ML-based software vulnerability patch data curation process aims to address both informativeness and quality issues by measuring both epistemic and aleatoric uncertainties.This dual uncertainty measurement, along with our EHAL heuristic, defined in Section IV, aims to enhance the efficacy of ML-based vulnerability data curation.
@@
Researchers in engineering and scientific domains increasingly use uncertainty quantification to address measurement errors and model uncertainties [43, 44]. While UQ studies provide numerous approaches to measure uncertainty, they also reveal that specific datasets and data types introduce domain-specific challenges and there is no one-fits-all uncertainty estimate [45, 46]. For instance, recent studies emphasize the presence of evolving distributions, noise, and the need for domain-specific UQ methods tailored to structured representations and unique patterns in source code data [47, 48]. In this first RQ, we aim to identify the most suitable UQ approximation techniques for addressing the unique noise and challenges associated with software vulnerability data curation. Applying UQ to vulnerability datasets is particularly challenging because the domain of vulnerability patches and descriptions is unique and therefore it is not possible to determine in advance which UQ method will perform best. Vulnerability patches listed in public datasets vary significantly in quality and utility, necessitating precise UQ measurements to account for these variations.
There are numerous approaches for quantifying epistemic and aleatoric uncertainty. Our objective is to identify suitable UQ approaches for patch data curation, focusing on neural network-based models, as they dominate ML-based curation tasks. We prioritize methods that integrate seamlessly without architectural modifications and can handle data distribution shifts common in vulnerability patches [48].
Probabilistic UQ methods are categorized into Bayesian and Frequentist approaches [39]. While conformal prediction, a popular Frequentist method, assumes data exchangeability and offers only marginal coverage guarantees [49, 50], Bayesian approaches provide more flexible uncertainty estimation. However, full Bayesian networks are computationally prohibitive. Instead, we adopt Bayesian approximations, specifically Monte Carlo Dropout and Variational Inference, which deliver instance-level uncertainty efficiently and integrate well with existing models.
To answer the question of which UQ approaches are suitable for the specific domain of vulnerability patch datasets, we first introduce three UQ approximation techniques in Section III-A. Following that, Section III-B discusses two different assumptions underpinning the representation of vulnerability data, specifically whether it is homoscedastic or heteroscedastic. Subsequently, we examine metrics that can effectively gauge the accuracy of the measured uncertainty in Section III-C. Finally, we discuss the evaluation experiment setup and the results in Section III-D and Section III-E, respectively.
@@ III-A2 Monte Carlo Dropout
Monte Carlo Dropout is a popular regularization technique used in training neural networks, where a neural network layer’s output is connected to a dropout layer. With it, we drop, i.e, zero out, the activation of the connected neurons at probability $P_{d}$ during training. When dropouts are used for UQ, we also turn on dropout during inference time, which provides a sampling of a single set of neural network weights [16]. In essence, with Monte Carlo Dropout, we approximate probability distribution $p(y|x,D,\theta)$ with a discrete probability distribution, i.e., $p(y|x,D,\theta_{i})$, $i=0,1,2,\ldots,T-1$ by sampling multiple sets of neural network weights with $T$ stochastic passes.
@@ III-C Uncertainty Measures
There is often a need to disentangle aleatoric and epistemic uncertainties [58, 39], as this work also demonstrates. The Law of Total Variance states that total variance can be decomposed into two components,
$Var(y)=E[Var[y|x]]+Var[E[y|x]]$ (4)
where, in this study, $x$ represents a single commit patch and $y$ its vulnerability status. A patch identification model, a neural network, approximates the function $y=\phi(x)$. We treat $Var(y)$ as a representation of total uncertainty. Since $E[y|x]$ is expected to average out the noise, the variance of $E[y|x]$ can be attributed to the model’s lack of knowledge during inference. As such, we consider $Var[E[y|x]]$ an estimate of epistemic uncertainty. Subtracting the epistemic component leaves the aleatoric uncertainty. Accordingly, we regard $E[Var[y|x]]$ as an estimate of aleatoric uncertainty. Given this understanding, in the context of homoscedastic models, the literature typically refers to the entropy defined in equation (1) as a measure of total uncertainty, the quality defined in equation (III-C) as a measure of data or aleatoric uncertainty, and the mutual information estimated in equation (2) as a measure of model or epistemic uncertainty [58, 59, 60]. In the context of heteroscedastic models, we obtain two probabilistic distributions, one representing aleatoric uncertainty, $p_{ale}(y|x)$, estimated from the expectation of $\sigma$, and the other representing epistemic uncertainty, $p_{epi}(y|x)$, computed using the variance of $\mu$ (see Figure 3). With these two distributions, we can compute two entropy values, denoted as $\mathcal{H}_{ale}(y|x)$ and $\mathcal{H}_{epi}(y|x)$ according to equation (1), which correspond to aleatoric and epistemic uncertainties, respectively. A detailed explanation of the application of the Law of Total Variance to separate total uncertainty into epistemic and aleatoric uncertainty can be found in the machine learning literature [58, 39, 20].
@@ III-D2 Datasets and Feature Representation
III-D3 Evaluation Procedures
We implement six deep learning models by combining three UQ techniques (Section III-A) with two model architectures (Section III-B). First, we develop Vanilla models for both homoscedastic and heteroscedastic noise assumptions (Figures 2 and 3). Each model uses CodeBERT for feature extraction, followed by three fully connected layers with dropout. The homoscedastic model applies softmax, while the heteroscedastic model outputs $\mu$ via ReLU and $\sigma$ via softplus to ensure positivity. We extend these base models to support Monte Carlo Dropout by enabling stochastic dropout during training and inference, and Model Ensemble by training five independent models and averaging their predictions.
We use the Adam optimizer during training. For homoscedastic models, we train with the negative log likelihood loss function (i.e., mathematically, the cross-entropy loss in two-class classification), and for heteroscedastic models, we train with a stochastic negative log likelihood loss function introduced by Kendall and Gal [17].
We tune hyperparameters using a validation dataset and monitor training with early stopping, halting after five epochs of no improvement in validation loss. Each layer contains 300 neurons with a dropout rate of 0.1. The model checkpoint with the lowest validation loss is selected for evaluation.
To answer RQ1, we design and carry out a series of numerical experiments. For each experiment, we follow standard training, validation and testing protocols. Unless otherwise stated, the ratio of training and test datasets is $0.8:0.2$. From the training dataset, 10% is set aside as a validation dataset. During training, the majority class, i.e., non-vulnerability patches, are under-sampled to have a balanced training dataset, while during evaluation the test dataset is kept as is.
@@ III-E2 Separating Epistemic from Aleatoric Uncertainty
In this section, we aim to ascertain whether the best UQ model we evaluated above, i.e., the heteroscedastic Model Ensemble, can serve as a reliable base for accurately outputting the disentangled uncertainties, i.e., the epistemic and aleatoric uncertainties.We demonstrate that epistemic and aleatoric uncertainties behave as expected when curating patch data. In Table III, we set aside 80% of patches as training data, train the patch identification model with increasing training dataset sizes (60%, 80%, and 100% of the training data), and then estimate the epistemic and aleatoric uncertainties on the test dataset (the remaining 20% of the patches). The table shows that as we increase the training data size, there is a significant reduction in epistemic uncertainty, i.e., a 1.5% to 9.3% decrease in epistemic uncertainty when the training data increased from 60% to 80% and then to 100%. In contrast, the aleatoric uncertainties saw smaller changes. Since an increase in the training data size indicates an improvement in the model’s knowledge about the problem space, which is captured by a reduction in epistemic uncertainty, it follows that for data curation of software vulnerability patches, a UQ technique that disentangles epistemic and aleatoric uncertainties is required.
@@
Data $\overline{\mathcal{H}}_{epi}$ $\Delta_{epi}\%$ $\overline{\mathcal{H}}_{ale}$ $\Delta_{ale}\%$ 60% 0.794 0.0% 0.817 0.0% 80% 0.782 $\frac{0.794-0.782}{0.794}\approx 1.5\%$ 0.819 $\frac{0.817-0.819}{0.817}\approx-0.2\%$ 100% 0.709 $\frac{0.782-0.709}{0.782}\approx 9.3\%$ 0.790 $\frac{0.819-0.790}{0.819}\approx 3.5\%$
@@
Finally, we pose the question of whether modeling the noise in software patch data as heteroscedastic offer any advantage over modeling it as homoscedastic. To optimize computational resources, we focused on the heteroscedastic Model Ensemble, the best-performing UQ approximation, for comparison against homoscedastic UQ techniques.The results in Table IV are obtained by averaging over all data quality shift intensities (from 0 to 0.4 with a step size of 0.1). The results show that the heteroscedastic model offers a significant advantage in predictive performance (higher F1-score) and superior or similar UQ quality (lower or nearly equal Brier Score). Given this observation, we conclude that for vulnerability patches, the UQ measures offered by the heteroscedastic models are more accurate than the homoscedastic models, due likely to the fact that the data distribution in vulnerability patches can vary between projects (and within a project over time), which is more accurately captured by heteroscedastic models.
@@
1. From the three UQ models we considered, Model Ensemble produces the most accurate UQ estimation in response to artificial data quality degradation. 2. To effectively assess the knowledge gained by a machine learning-based automated patch curation, it is crucial to separate epistemic and aleatoric uncertainty. Epistemic and aleatoric uncertainty are distinguishable during automated patch data curation. 3. Heteroscedastic models provide more accurate UQ estimation for patch data curation.
Based on RQ1, we design and evaluate the data curation system. First, we assess the algorithm’s ability to predict patch vulnerability. Second, we examine its impact on vulnerability prediction, a key downstream application (Figure 5).
We begin with a small set of high-quality seed patches $(X_{s},Y_{s})$ to train a UQ model. Next, we curate patches from a pool $X_{p}$ of potentially low-quality or project-specific patches. To prevent overlap, we ensure $X_{p}\cap X_{s}=\emptyset$. The curation process involves two steps: running inference on $X_{p}$ and applying the EHAL heuristic to select informative, high-quality patches.
[…]
Figure 5: Evaluation setting for RQ2.
IV-A Automatic Curation of Vulnerability Patches with UQ
Given an accurate estimation of epistemic and aleatoric uncertainty for patch data curation, as described in RQ1, it is still an open problem of how we use these two quantities to improve patch data curation. In this RQ2, we focus on the practical aspects of using UQ in ML-based patch data curation in order to obtain higher quality datasets. To this end, we design the EHAL Heuristic (Algorithm 1) that actively and selectively adds patches based primarily on high epistemic uncertainty to include in the dataset. The algorithm relies on a pre-trained UQ model, which takes as input a pool of commit patches and outputs a selected subset. These input patches may be either labeled or unlabeled. For unlabeled patches, the algorithm identifies candidates for annotation, while for labeled ones, it selects instances that are both informative and reliable for training.
Instead of leveraging a formula that combines both epistemic and aleatoric uncertainties, the EHAL heuristic takes a “select-and-then-reject” approach. We opted for this strategy over scalar combinations (e.g., weighted sums of uncertainties) to avoid arbitrary tradeoff calibration. Empirically, we found this approach simpler to tune and less sensitive to outliers in either uncertainty dimension. As shown in Algorithm 1, it first identifies patch instances with the highest epistemic uncertainty as candidates, but rejects them if they are among the patches with the highest aleatoric uncertainty, i.e., it rejects those on the lowest extreme of the quality spectrum. The name of the heuristic is drawn from this idea as well: Epistemic High Aleatoric Low (EHAL).
More specifically, the algorithm’s main function, DataCurationWithEHAL (Lines 1–8), curates $n$ patches from a candidate pool $\mathbb{X}_{\text{pool}}$, guided by their pre-computed epistemic and aleatoric uncertainties, $\mathbb{U}_{\text{pool}}$. It iteratively selects patches by invoking the EHAL heuristic (Lines 9–19), which forms the core of the algorithm. The heuristic identifies the candidate patch with the highest epistemic uncertainty using the TopOneByEpistemicUQ operation (Lines 12 and 17). Simultaneously, the top $n_{\text{ale}}$ patches with the highest aleatoric uncertainty are determined using TopNByAleatoricUQ (Lines 13 and 18), where $n_{\text{ale}}$ is a hyperparameter determined empirically. If the candidate with the highest epistemic uncertainty is also among the top aleatoric candidates, it is excluded from consideration, and the selection process iterates. This ensures that the final curated dataset includes only patches that are both highly informative and reliable, effectively balancing the trade-off between informativeness (epistemic uncertainty) and quality (low aleatoric uncertainty).
To evaluate the EHAL Heuristic, we consider two baselines. We construct a data selection heuristic completely opposite to the EHAL Heuristic, i.e., Epistemic Low Aleatoric High (ELAH). ELAH selects the patches with the lowest epistemic uncertainty, and rejects them if they are also among those with the lowest aleatoric uncertainty. In addition, we also compare our algorithm with the random baseline that selects patches randomly. Random baselines are frequently used in the active learning literature [63]. By comparing with a random baseline, we can determine whether the data curation algorithm has any knowledge about data selection at all.
In our evaluation, we gradually expand the training dataset, using one of the three algorithms: the proposed one and the two baseline algorithms. For each, we divide the training data partition randomly into two parts: the initial training dataset (20% of available patches) and the candidate training instances pool (60% of available patches). We construct the training dataset, starting with the initial training dataset, and train a patch identification model. With the trained model, we estimate uncertainty measures, such as aleatoric and epistemic uncertainties. Based on the uncertainties, we select 10% of instances from the candidate pool and add these to the training dataset. We then retrain the model and repeat the process. We stop when we exhaust the candidate dataset. To evaluate the model’s predictive performance, we compute the F1 score on the test dataset (20% of available patches) as the size of the training dataset increases. To ensure the reliability of the results, we repeat the process 10 times, each beginning with a random shuffling of the entire dataset.
Input: $\mathbb{X}_{\text{pool}}$: Pool of candidate patches
$\mathbb{U}_{\text{pool}}$: Uncertainty values for candidate patches
$n$: Number of patches to curate
$n_{\text{ale}}$: Hyperparameter for the number of aleatoric instances to consider in each iteration
Output: $\mathbb{D}_{\text{new}}$: Final list of curated patches